The Vertical Wien filter HV was ramped up yesterday to 20.5 kV (Digital voltage readout is 1.5 % lower than set point. For 100 deg spin rotation at 200 keV the required voltage per plate is ~20.2 kV ) and was left soaking overnight. There were no vacuum spikes during the soak and the voltage did not trip off.
I have ramped the HV down to 0 V in preparation for Horizontal Wien filter HV evaluation.
Day #159, Week 32 Day 4, Wrap-Up:
* TOSP and THA for beam operations in Injector without
NL PSS certification complete released -- must read for
Injector beam commissioning shift-takers and those
involved with upcoming CEBAF PSS Certification.
* PCup, FC1, and FC2 ItoV HCO completed. 3876865
* 500keV Dump picoammeter HCO completed, completing
Injector picoammeter system HCO. 3876873
* Beam sync verified at InjSB ItoV chassis. Beam Sync
and 30Hz sync signals verified on MCC scope. No issues.
Shukui and I were very excited to measure photocathode QE in Gun2.
We're using laser bypass key (see fig. 1) but TMG shutters won't open....seems like 18V maybe not getting there? b/c none of the lasers will pass light. On the other hand, the PSS shutters go green, that's a good thing.
Anyway, no light to the photocathode means no QE scan yet, no PSS laser shutter test, etc.
TS'ing under way!
Gun valve solenoid would not close gun valve today, so I hard-wired air line directly to valve, closed fine.
Air blows strong through solenoid exhaust, rather than into valve mechanism.
Same old style valve.
Same air pressure (right?!?)
On Phil's punchlist for Friday...and then maybe a QE scan, so maybe window for NEG activations has passed, photocathode in the gun, maybe making beam on Monday, checking QE issues, etc, blah blah.
Vacuum in the 1L injector beam line was monitored while the Vertical Wien filter voltage was ramped simulatenously for the Pos and the Neg power supplies. First in the Positive Wien angle configuration, then in the Negative Wien angle configuration.
Figure 1 shows the vacuum response of the ion pumps shown in Figure 2 to high voltage.
In the positive Wien angle configuration, the voltage was ramped in 1 kV stpes about every 2-3 minutes. At 11 kV a vacuum spiked occurred. At 19.5 kV, the HV enable permit was lost in both Pos and Neg power supplies.
The purpose of the attached TOSP (ACC-21-115759-TOSP) is to define the controls, procedure and responsibilities
for safely running beam to the 500 keV region of the injector without the possibility of sending beyond that point
before and interleaved with the weeks-long process of PSS certification.
Those who stand shift (OPS, CIS, CASA) to run beam and SSG staff who carry out the PSS Certs must read,
understand and sign the TOSP which will be posted in the MCC.
Later today, the Injector will be swept and locked up for PSS certifications. Lock-up will occur after 430p, allowing work in the tunnel wrap-up for the afternoon.
Thank you DC Power for enabling operation of MBO0I06 for another night, mid-HCO.
IHA0I02 - stepper motor replaced, stepper motor control module replaced, linear pot replaced, preamp amplifier replaced, reterminated the control cables and read back cables, cycled harp and verified the stepper motor and limit switch read backs are operational (*ITV0I02 control voltage is disabled), still need to test the anti-collision circuit that prevents the viewer from inserting while the harp is "IN"
*Need to speak with Joe and Phil about interlocking the harp with the viewer (i.e.
PWF1I04 - used an separate cable to run the HV interlock for the vertical Wien bottom, reterminated both inter lock cables, reconfigured the pin-out on the interlock boxes on the cable tray, verified the power supply interlock "opens" when the field cables are disconnected.
PWF1I06 - used a new cable to run both interlocks for the left and right HV field cables, terminated cables and reconfigured the pin-out on the interlock boxes located on the cable tray, replaced the 24V relay inside the switcher chassis for the negative side power supply, verified the power supply interlock "opens" when t